OpenStudy (anonymous):

Find the explicit function for this recursive equation: \[ y = 2y+5 \] Is this possible? What do you guys think?

OpenStudy (anonymous):

\[ y = 2(2y+5)+5 = 4y + 15 \]

OpenStudy (anonymous):

One solution is \(y=-5\). A constant function

OpenStudy (anonymous):

I think that's the only solution :/

OpenStudy (anonymous):

How about just \[ y=2y \]

OpenStudy (anonymous):

\(y\) would have to be 0 for that to be true.

OpenStudy (anonymous):

No functions have this property then

OpenStudy (anonymous):

other than 0

OpenStudy (anonymous):

Well, constant functions are still functions, just not very interesting ones.

OpenStudy (anonymous):

I'm thinking maybe some periodic function might work, but then you'd have to restrict the domain.

OpenStudy (anonymous):

How about something like: \[ y(x-1) = 2 y(x) +5 \]

OpenStudy (anonymous):

\[ y(x-2) = 2(2y(x)+5)+5 \]

OpenStudy (anonymous):

\[ y(x-n) = 2^ny(x) + \sum_{k=1}^{n}2^{k-1}(5) \]

OpenStudy (anonymous):

I kinda wonder how something like this might be solved.

OpenStudy (anonymous):

By the way, did you manage to get quoting to work?

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio By the way, did you manage to get quoting to work? \(\color{blue}{\text{End of Quote}}\) Indeed I did.

OpenStudy (anonymous):

Anyway, I don't know if there's a rigorous way to prove it, but I think that, generally, a function shifted horizontally by one unit can never be the same as the original function scaled by a factor of 2 and shifted vertically by five units.

OpenStudy (anonymous):

unless it's constant, like the first two you posted.

OpenStudy (anonymous):

**an appropriate constant**

OpenStudy (anonymous):

Well, if you seeded it like said... \(y(0) = 0\), would the rest of it sort of fall into play?

OpenStudy (anonymous):

\[ y(x-1) = 2y(x) +5 \]So \[ y(x) = \frac{y(x-1)-5}{2} \]And \[ y(1) = \frac{y(0)-5}{2}=\frac {-5}2 \]

OpenStudy (anonymous):

You'd be able to list out a bunch of inputs and values, so it would be a discrete function of sorts, so shouldn't there be some way to describe this function, even if not algebraically?

OpenStudy (anonymous):

Ah, right, you could express every output in terms of \(y(0)\)... I get the feeling this would give you a linear function.

OpenStudy (anonymous):

Or maybe not: In this case, \(y(x)=-\dfrac{5}{2}x\), but \(y(x-1)\not=2y(x)+5\) when you plug it in.

OpenStudy (anonymous):

\[ y(2) = \frac{\frac{-5}2-5}{2} = \frac{\frac{-15}2}{2} = \frac{-15}4 \]

OpenStudy (anonymous):

You could find a pattern if you kept doing this.

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio \[ y(x-n) = 2^ny(x) + \sum_{k=1}^{n}2^{k-1}(5) \] \(\color{blue}{\text{End of Quote}}\) Inverting this:\[ y(x+n) = \frac{y(x)- 5\sum_{k=1}^{n}2^{k-1}}{2^n} \]Now letting \(x=0\) and \(n=x\):\[ y(0+x) = y(x) = \frac{y(0)- 5\sum_{k=1}^{x}2^{k-1}}{2^x} = \frac{- 5\sum_{k=1}^{x}2^{k-1}}{2^x} \]

OpenStudy (anonymous):

Since it is a geometric series, the summation can be simplified into an algebraic expression.

OpenStudy (anonymous):

But you're only dealing with integeres \(x\), right? A real-valued function would be difficult to describe in this way.

OpenStudy (anonymous):

Well, the recursive equation provides a definition to verify integer domain. Any real valued function which hits all the integer values correctly would be a valid extension onto the real domain.

OpenStudy (anonymous):

In the same way that \(\Gamma(n+1)\) extends \(n!\)

OpenStudy (anonymous):

Hmm, can't say I know enough about this stuff to be able to contribute here, but it's interesting nonetheless.